what is energy of a photon that has a wavelength of 535nm? Show work.
Answer #1


Given:
lambda = 535 nm = 5.35*10^-7 m

use:
E = h*c/lambda
=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(5.35*10^-7 m)
= 3.716*10^-19 J
Answer: 3.72*10^-19 J